A 0.752-g sample of an unknown gas has a volume of 530mL and a pressure of 622 mmHg at 45.1°C. calculate the molar mass of this compound.
Answer #1

From ideal gas equation

PV = nRT

PV = (m/M) ×RT

Where, m = mass

M = molar mass

(622/760) × 0.530 =( 0.752/M)×0.0821×318.1

M =45.3 g/mol = 45.0 g/mol